OpenStudy (anonymous):

For the following system, use the second equation to make a substitution for y in the first equation. 3x + y = 2 y + 4 = 3x

OpenStudy (anonymous):

3x+ 3 - 4x= 2?

OpenStudy (anonymous):

solve for y on the second equation

OpenStudy (anonymous):

-1?

OpenStudy (anonymous):

y + 4 = 3x y + 4-4 = 3x-4 y = 3x-4 Plug that into 3x + y = 2 What do you get then?

OpenStudy (anonymous):

3x+ 3x- 4 = 2?

OpenStudy (anonymous):

Yes you are correct. Then solve for x